# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ A.2 Proof of Theorems 1, 2 and 3
Applying Lemma A.7 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$ gives the desired result.
Appendix B Details of Numerical Experiments
All experiments are conducted on a Slurm cluster. For experiments with 4 GPUs, each node has either 4 NVIDIA GEFORCE RTX 2080 Ti (11GB) GPUs or 4 NVIDIA A40 (48GB) GPUs or 4 NVIDIA L40s (48GB) GPUs. For experiments with 8 GPUs, each node has 8 NVIDIA A100 (80GB) GPUs. For CPUs, each node has either 64 cores or 96 cores. The total RAM is 512GB or 1TB.
B.1 Three-layer Convolutional Neural Network on MNIST
Table 4: Training hyperparameters for three-layer CNN on MNIST
Model Three-layer CNN on MNIST
Training samples 6M (100 epochs)
Weight initialization Default
Optimizer Adam
Learning rate schedule None
$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
$\varepsilon$ $10^{-8}$
Base learning rate 0.0008
Base batch size 8
Maximum batch size 50000
Weight decay 0
Base global batch size 4096
Base local batch size 1024
Maximum global batch size 60000
Maximum local batch size 15000
Weight decay 0
Figure 6: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies with three different seeds.
Figure 7: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
Figure 8: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
Figure 9: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
B.2 ResNet-50 on CIFAR-10
@@
Training samples 10M (200 epochs)
@@
Learning rate warmup (samples) 1M (10%)
@@
Minimum learning rate 0.005
Base learning rate 0.05
@@
Maximum global batch size 50000
Maximum local batch size 12500
Weight decay 0.0005
@@
Figure 10: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with three different seeds.
Figure 11: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
Figure 12: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
Figure 13: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
@@
Figure 14: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with 8 workers with three different seeds.
@@
Figure 15: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with 8 workers with two other different seeds.
@@
Figure 16: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with 8 workers with three different seeds.
@@
Figure 17: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with 8 workers with three different seeds.

### v2

@@ B.2 Proofs of Theorems 1, 2 and 3
Applying Lemma B.7 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$ gives the desired result.
Appendix C Details of Numerical Experiments
For experiments with 4 GPUs, each node has either NVIDIA L40S (48GB; for ResNet-50 on CIFAR-10) or A40 (48GB; for MicroLlama 300M on C4) or A100 (SXM4 80GB; for ResNet-101 on ImageNet) GPUs.
C.1 ResNet-50 on CIFAR-10
@@
Training samples 30,000,000 (600 epochs)
@@
Learning rate warmup (samples) 1,000,000 (10%)
@@
Base learning rate 0.005
Learning rate scaling rule linear (for constant batch sizes)
@@
Maximum global batch size 50,000
Maximum local batch size 12,500
Weight decay 0.0001
@@
Precision bfloat16
Data-parallel size 4
(a) $H=1$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
@@
Figure 3: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10.
Table 4: Adaptive batch size strategies for ResNet-50 on CIFAR-10; mean (standard deviation) of no. of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best validation accuracy (acc.; in %) of all three seeds
Schedule $H=32$ $H=16$
steps time bsz. acc. steps time bsz. acc.
Constant 1824 0.98 (0.00) 4096 75.23 (7.19) 1824 0.99 (0.00) 4096 79.62 (4.39)
Constant 896 0.95 (0.00) 8192 33.92 (10.78) 912 0.98 (0.00) 8192 31.23 (14.77)
Constant 576 1.07 (0.00) 12500 10.70 (0.79) 592 1.10 (0.00) 12500 14.92 (5.49)
$\eta=0.8$ 950 (98) 1.14 (0.02) 7787 (607) 71.49 (3.09) 827 (25) 1.14 (0.01) 9030 (209) 72.15 (3.82)
$\eta=0.85$ 1088 (0) 1.16 (0.02) 6943 (129) 70.54 (1.03) 880 (43) 1.14 (0.01) 8457 (374) 72.41 (2.52)
$\eta=0.9$ 1163 (67) 1.14 (0.01) 6346 (341) 73.02 (2.82) 1072 (43) 1.15 (0.00) 7027 (238) 73.22 (3.72)
Schedule $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc.
Constant 1828 1.06 (0.00) 4096 88.72 (0.52) 1831 1.34 (0.00) 4096 89.31 (0.34)
Constant 912 1.01 (0.00) 8192 76.56 (2.29) 915 1.15 (0.00) 8192 78.19 (2.63)
Constant 596 1.14 (0.00) 12500 35.70 (11.59) 599 1.23 (0.00) 12500 40.60 (15.22)
$\eta=0.8$ 750 (25) 1.16 (0.00) 10001 (346) 74.66 (0.88) 1023 (211) 1.36 (0.05) 7548 (1554) 80.56 (2.08)
$\eta=0.85$ 771 (41) 1.16 (0.01) 9736 (494) 75.57 (0.21) 935 (297) 1.35 (0.08) 8525 (2399) 78.89 (4.03)
$\eta=0.9$ 756 (14) 1.16 (0.01) 9916 (166) 74.77 (0.42) 1246 (354) 1.42 (0.07) 6415 (2094) 81.73 (3.65)
Observations.
Note that due to the stochastic nature of drawing data samples when computing gradients, the adaptive batch size strategies might give rise to different adaptive batch size schedules when different random seeds. We hence perform the training runs for three different random seeds to evaluate the robustness of the proposed strategies (see results of the other two seeds in Figures 4 and 5). We observe from Table 4 that, as the number of local steps $H$ grows, constant batch sizes might have more varied performance (in terms of standard deviation of best validation accuracy) than the proposed strategies.
Also notice that, due to the use of relatively large batch sizes, all the experiments make use of the technique of gradient accumulation, which essentially consists of serial operations using for-loops. Therefore, the benefit of using large batch sizes is not reflected in the total training time. If a larger number of GPUs or stronger GPUs with more memory are used, training with large batch sizes will consume substantially shorter time. Consequently, it is better to measure the computational efficiency of different training runs using the number of total training steps (Shallue et al., 2019).
(a) $H=32$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(b) $H=16$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(c) $H=4$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(d) $H=1$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(e) Adaptive batch size strategies for $H\in\{32,16,4\}$
@@
Figure 4: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10 of the second seed.
(a) $H=32$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(b) $H=16$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(c) $H=4$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(d) $H=1$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
(e) Adaptive batch size strategies for $H\in\{32,16,4\}$
@@
Figure 5: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10 of the third seed.
C.2 MicroLlama 300M on C4
Table 5: Training hyperparameters for MicroLlama 300M on C4
Model MicroLlama 300M on C4
Training samples (sequences) 2M
Sequence length 2048 tokens
Weight initialization Default
Normalization Layer Normalization
Optimizer AdamW
$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
Learning rate schedule Linear warmup + cosine decay
Learning rate warmup (samples) 20,000 (1%)
Peak learning rate 0.001
Base learning rate 0.0001
Learning rate scaling rule linear (for constant batch sizes)
Base global batch size 256
Base local batch size 64
Maximum global batch size 8,192
Maximum local batch size 2,048
Weight decay 0.1
Weight decay skip bias Yes
Precision bfloat16
Dropout 00
Gradient clipping $1.0$
Data-parallel size 4
Table 6: Adaptive batch size strategies for MicroLlama 300M on C4; mean (standard deviation) of no. of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best validation loss (cross-entropy loss; estimated by 100 iterations) of all three seeds
Schedule $H=32$ $H=16$
steps time bsz. loss steps time bsz. loss
Constant 31744 10.55 (0.04) 512 4.20 (0.18) 16384 10.51 (0.03) 512 5.88 (0.21)
Constant 16384 10.51 (0.03) 1024 4.96 (0.21) 7936 10.60 (0.60) 1024 5.08 (0.21)
Constant 8192 9.77 (0.01) 2048 5.88 (0.21) 4096 10.47 (0.02) 2048 5.84 (0.10)
$\eta=0.8$ 14336 (1774) 11.00 (0.12) 1162 (155) 4.81 (0.31) 6059 (391) 11.02 (0.05) 1356 (89) 4.96 (0.18)
$\eta=0.9$ 13995 (2132) 11.15 (0.44) 1207 (134) 4.88 (0.19) 7168 (678) 11.11 (0.36) 1163 (130) 4.72 (0.19)
Schedule $H=4$
steps time bsz. loss
Constant 14336 11.00 (0.12) 512 4.81 (0.31)
Constant 1968 11.28 (0.04) 1024 5.14 (0.13)
Constant 992 10.92 (0.03) 2048 5.99 (0.04)
$\eta=0.8$ 1206 (177) 11.14 (0.02) 1701 (260) 5.15 (0.38)
$\eta=0.9$ 1238 (199) 11.14 (0.06) 1668 (303) 4.94 (0.33)
Observations.
For the task of language modeling with MicroLlama 300M on the C4 dataset, we also perform the training runs with three different random seeds. Similar conclusions to those in the main text can also be made, even though the proposed adaptive batch size strategies lead to different batch size growth patterns when using different random seeds (see Figures 6 and 7).
(a) $H=32$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
(b) $H=16$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
(c) $H=4$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
(d) Adaptive batch size strategies for $H\in\{32,16,4\}$
Figure 6: Validation loss and local batch sizes of Local AdamW with adaptive batch size strategies for MicroLlama 300M on C4 of the second seed.
(a) $H=32$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
(b) $H=16$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
(c) $H=4$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
(d) Adaptive batch size strategies for $H\in\{32,16,4\}$
Figure 7: Validation loss and local batch sizes of Local AdamW with adaptive batch size strategies for MicroLlama 300M on C4 of the third seed.
C.3 ResNet-101 on ImageNet
We also train a ResNet-101 (He et al., 2016) on the ImageNet dataset (Russakovsky et al., 2015) to evaluate our method on models and datasets of larger scales.
Table 7: Training hyperparameters for ResNet-101 on ImageNet
Model ResNet-101 on ImageNet
Training samples 256,233,400 (200 epochs)
Weight initialization Default
Optimizer Momentum SGD (SHB)
Learning rate schedule Linear warmup + cosine decay
Learning rate warmup (samples) 6,405,835 (2.5%)
Peak learning rate 2.5
Base learning rate 0.25
Base global batch size 512
Base local batch size 128
Maximum global batch size 52,000
Maximum local batch size 13,000
Weight decay 0.0001
Momentum 0.9
Precision bfloat16
Data-parallel size 4
Table 8: Adaptive batch size strategies for ResNet-101 on ImageNet; no. of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best top-1 validation accuracy (acc.; in %), and best top-5 validation accuracy (acc.@5; in %)
Schedule $H=32$ $H=16$
steps time bsz. acc. acc.@5 steps time bsz. acc. acc.@5
Constant 10656 14.56 6000 59.20 81.84 10672 14.78 6000 63.76 85.18
Constant 4896 14.35 13000 38.77 63.30 4912 14.34 13000 50.87 74.89
$\eta=0.9$ 5216 14.53 12284 50.61 74.59 5072 14.64 12603 55.63 78.86
$\eta=0.95$ 5280 14.31 12124 49.13 73.23 5088 15.09 12573 58.41 81.17
Schedule $H=4$
steps time bsz. acc. acc.@5
Constant 10676 17.20 6000 71.28 89.97
Constant 4924 15.41 13000 62.66 84.33
$\eta=0.9$ 4952 15.62 12931 65.90 86.47
$\eta=0.95$ 4976 16.75 12873 67.05 87.24
Observations.
We observe from Table 8 that the proposed adaptive batch size strategy with $\eta=0.9$ achieves better generalization performance (in terms of both accuracy and top-5 accuracy) than that of a constant large local batch size of 13,000 but worse than that of a constant small local batch size of 6,000 (i.e., generalization gap), while requiring also an intermediate amount of training time. The performance gap grows with the number of local steps $H$ (cf. Figures 8, 9 and 10).
@@ C.3 ResNet-101 on ImageNet
Figure 8: Validation accuracy, top-5 validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies for ResNet-101 on ImageNet for $H=32$.
@@
Figure 9: Validation accuracy, top-5 validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies for ResNet-101 on ImageNet for $H=16$.
Figure 10: Validation accuracy, top-5 validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies for ResNet-101 on ImageNet for $H=4$.